OpenStudy (anonymous):

The hillside Garden Club wants to double the area of its rectangular display of roses. If it is now 6 meters by 4 meters, by what equal amount must each dimension be increased?

OpenStudy (anonymous):

6*4=24 24*2=48 trail and error 8*6=48 answer 2..